A light ray travelling from glass to air undergoes refraction. Calculate the angle of refraction to the nearest whole number if the incidence angle is 30° and the refractive for air and glass are 1.0 and 1.5 respectively.

1
Expert's answer
2022-02-09T09:14:50-0500

Answer

Refractive index are

n=1.5

n'=1.0

Incidence angle "\theta=30°"

Using snell law

So angle of refraction

"\theta'=sin^{-1}( \frac{n sin \theta}{n'}) \\=sin^{-1}( \frac{1.5\times sin 30°}{1}) \\=48.59°"